# Defining Boundaries: A Spectrum of Task Feasibility for Large Language Models

### v1

@@ 3.1 Evaluation Setup
Metrics. We evaluate distinguishability using two metrics: the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Kolmogorov–Smirnov Statistic (KSS). The AUROC measures the probability that a model ranks a randomly selected positive instance higher than a randomly selected negative instance. An AUROC value of 1.0 signifies perfect classification accuracy, whereas a value of 0.5 indicates no better performance than random guessing. The KSS assesses the maximum distance between the cumulative distribution functions of two sets of samples, with higher values indicating greater separation between distributions. In addition, we assess model calibration, which examines the correspondence between a model’s expressed confidence and its actual accuracy. "We selected the Brier Score as our metric because it favors probability predictions that are both well-calibrated and precise, aligning with our purpose.
Table 4: GPT-4 can still distinguish feasible and unfeasible data when the dataset generation was switched to GPT-3.5.
Method AUROC KSS Brier Score
Pre 0.941 0.805 0.091
3.2 Results and Analyses
Table 2 presents the outcomes of various methods used to derive confidence scores from different LLMs. We provide a summary of several critical insights from these experiments. $\mathbf{1}$. Excluding GPT-3.5, the pre-response method generally outperforms other models. This suggests that adding additional thinking steps does not typically enhance performance for LLMs. This finding is particularly notable, as in prior studies—like differentiating between ambiguous and unambiguous questions (Hou et al., 2023) or identifying known versus unknown queries (Liu et al., 2024)—increased reasoning steps (e.g., mid, post, and mix-response) proved advantageous. A plausible explanation is that for more advanced LLMs, the process of identifying feasible versus unfeasible instructions becomes more straightforward. $\mathbf{2}$. Across all models and techniques, GPT-4 consistently delivers the most precise and well-calibrated confidence estimates through direct verbalization compared to other models, which is also shown in Fig. 3. Additionally, GPT-4 exhibits minimal variability in results across different methods; for instance, the AUROCs for pre and post are 0.965 and 0.967, respectively.
@@ 3.2 Results and Analyses
To further validate our results on more complex and real scenarios, we also create an additional benchmark dataset focused on long instructions, where each instruction comprises multiple tasks. The results in Table 3 indicate that long-form instructions are more challenging for current LLMs to accurately determine their feasibility compared to short-form benchmarks. For instance, GPT-4 using the pre-method achieved an AUROC of only $0.865$, significantly lower than the $0.965$ achieved in the previous short-form benchmark. Also, the overall calibration of probability becomes less well-aligned, which might make the model outputs less trustworthy. Those results highlight the increased difficulty of processing long-form instructions.
Our benchmark dataset was initially generated using GPT-4. To assess the potential data leakage, we conducted an ablation study where the dataset generation was switched to GPT-3.5. Subsequently, we evaluated GPT-4 using this newly generated dataset ($n=400$). The results are presented in Table 4 and show that GPT-4 can robustly and consistently distinguish feasible and unfeasible data.
4 Can We Teach LLM to Refuse Unfeasible Tasks without Hints?
With the benchmark, we observed that state-of-the-art LLMs can differentiate between feasible and infeasible tasks when provided with carefully designed query prompts. However, in practical applications, users typically interact with LLMs using straightforward queries without complex instructions. This raises a fundamental question: can we train LLMs to autonomously reject infeasible tasks during routine interactions without extensive prompting?
4.1 Methods
Given an instruction tuning dataset, we first reconstruct a refusal-added dataset where we explicitly incorporate refusal words into the data. We also find that some of the data belongs to infeasible tasks and might introduce hallucinations for refusal awareness. Here we have two strategies to achieve this.
4.1.1 Selection-based
To enhance the diversity of refusal expressions, we crafted multiple variations of refusal text. These expressions are detailed in Appendix D. For the identified infeasible data, we employ random sampling to select appropriate refusal expressions. This approach ensures a varied and comprehensive response strategy for handling queries that exceed the model’s capabilities.
4.1.2 Augment-based
Model Method OOD (Infeasible) Alpaca (Feasible)
OpenLLaMA-3b-v2 Original $0.105$ $0.357$ $\mathbf{0.059}$
Random $0.165$ $0.336$ $0.076$
Select $\mathbf{0.66}$ $0.335$ $0.086$
Augment $0.255$ $\mathbf{0.370}$ $0.069$
LLaMA2-7b Original $0.130$ $\mathbf{0.551}$ $0.070$
Random $0.140$ $0.296$ $0.184$
Select $\mathbf{0.735}$ $0.443$ $0.081$
Augment $0.175$ $0.432$ $\mathbf{0.065}$
LLaMA2-7b-chat $0.210$
GPT-3.5 $0.580$
GPT-4o $0.585$
4.1.3 Random-based
To underscore the significance of this selection process, we introduce an additional baseline, termed random-based, where queries are randomly chosen for response replacement. To ensure a fair comparison, we maintain the proportion of data undergoing response replacement consistent across different approaches.
Once the dataset has been augmented and structured, we proceed with standard supervised fine-tuning (SFT) on the newly constructed dataset.
4.2 Experimental Setting
Metrics. We assess the models from two dimensions: helpfulness and refusal awareness. To evaluate helpfulness, we leverage recent advancements in automated evaluation, using a high-performing large language model, specifically GPT-4o, as a proxy for human labeling. In this evaluation, the model ranks pairs of responses, one generated by the trained model and the other by a reference model. We use the average win rate as the metric for this assessment. To mitigate position bias, responses are presented in both sequential orders, and the average rank is calculated. The prompting template for evaluation is shown in Appendix E.
For evaluating refusal awareness, we implement lexical keyword matching to calculate the refusal rate. This method involves identifying specific keywords that signify abstention, apology, or denial, enabling us to measure the model’s capacity to appropriately refuse or defer a response when necessary.
Data. Alpaca dataset (Taori et al., 2023) is a widely used instruction dataset and we use its cleaned version as our main training dataset. We split the original dataset into training and testing. To evaluate helpfulness, we utilize the test part of Alpaca. To evaluate the ability of refusal, we utilize an OOD dataset from Sun et al. (2024). More summary statistics for the datasets we used can be found in Appendix B.
@@ 4.2 Experimental Setting
To more comprehensively measure the general ability of fine-tuned models, we tested the model on another dataset with a larger size called Alpagasus, which is mentioned in Chen et al. (2023). This dataset contains $700$+ data, carefully curated from multiple resources, and is regarded as "feasible" to LLMs. Since the models we fine-tuned were trained using the Alpaca dataset, we consider this scenario as an evaluation of their ability to handle out-of-distribution data.
4.3 Experimental Results
Trade-off between the helpfulness and refusal-awareness: We find this trade-off is similar to previous LLM studies (Bai et al., 2022; Touvron et al., 2023) when enhancing LLM’s instruction-following capabilities while ensuring they remain helpful and honest. We observe that there is a drop in general helpfulness. For example, in $3$b scale experiments, the win rate of select and random methods dropped nearly $2\%$ compared with original tuning (without refusal teaching). This is even worse with $7$b where all methods have over $10\%$ drop. This indicates that the proposed tunning methods still can’t provide a good balance between helpfulness and refusal-awareness.
Hard to improve general helpfulness: The results of testing the fine-tuned models on the Alpagasus dataset are shown in Table 9. The results show a general drop in win rate for all tunning methods compared with the original and suggest that these methods are not very resilient to distribution shifts and may not significantly improve general helpfulness. Therefore, future work should focus on developing more effective instruction-tuning methods to better manage distribution shifts.
5 Related Work
In this section, we review the progress on uncertainty quantification and hallucinations of large language models (LLMs).
5.1 Uncertainty Quantification in LLMs
Uncertainty quantification remains a core problem in deep learning. Guo et al. (2017) were among the first to point out that the predictive confidence of deep neural networks is often not well-calibrated. Recent studies have sought to address this by estimating and calibrating uncertainty specifically for language models (Xiao et al., 2022; Kuhn et al., 2023; Lin et al., 2023). A novel approach within this domain is verbalized confidence, which involves prompting LLMs to articulate their confidence levels in textual form Lin et al. (2022); Xiong et al. (2023). Tian et al. (2023) demonstrated that the method of verbalized confidence is effectively calibrated. Building on this straightforward approach, recent studies have further investigated its utility across various applications. These include tasks such as error detection (Xiao et al., 2022; Duan et al., 2023), ambiguity detection (Hou et al., 2023), and the identification of unanswerable queries (Liu et al., 2024). Our work can be seen as a generalization of utilizing the verbalized method in feasibility detection.
5.2 Hallucinations in LLMs
Despite the impressive performance characterized by high fluency and coherence, LLMs are still prone to generating unfaithful and nonfactual content, commonly referred to as hallucinations (Maynez et al., 2020). Several factors contribute to this phenomenon, including aspects of the training data, the algorithms used for training, and the inference processes (Ye et al., 2023; Zhang et al., 2023c; Rawte et al., 2023). Often, the training datasets themselves may include misinformation or become outdated, which can exacerbate the misalignment between the model’s outputs and factual accuracy (Penedo et al., 2023; Reddy et al., 2023; Li et al., 2024). Furthermore, LLMs have a tendency to overestimate their capabilities, leading them to produce incorrect responses with undue confidence and to struggle with recognizing when questions are unknown or unanswerable (Yin et al., 2023; Amayuelas et al., 2023; Cheng et al., 2024; Liu et al., 2024).
Recent research efforts have focused on addressing the issue of hallucinations in LLMs. For the detection of hallucinations, Azaria and Mitchell (2023) have developed a classifier that operates based on the internal states of LLMs. To measure the factuality of generated content, Lee et al. (2022) introduced a benchmark that utilizes both factual and nonfactual prompts. Furthermore, Varshney et al. (2023) employed an uncertainty-based approach to both detect and mitigate hallucinations during content generation. Zhang et al. (2023b) implemented a method that mimics human attention to factuality, guided by uncertainty scores. More recently, Sun et al. (2024) proposed out-of-distribution tasks but didn’t provide a formal definition and systematic summarization. There are also some recent recent works focused on investigating LLMs’ ability to abstain from answering to avoid hallucination (Slobodkin et al., 2023; Feng et al., 2024; Wen et al., 2024; Miyai et al., ). Our research contributes to this field by evaluating and training deliberate refusal of infeasible instructions, further aiding in the quantification and reduction of hallucinations in the era of LLMs.
6 Conclusion
We introduce the Infeasible Benchmark to analyze the behavior of LLMs when faced with instructions that exceed their capabilities. Our findings indicate that advanced LLMs are capable of distinguishing between feasible and infeasible tasks when provided with detailed guiding prompts. Yet, this capability diminishes in practical scenarios where users have minimal guidance regarding infeasible tasks. Additionally, we have developed fine-tuning methods aimed at enhancing the models’ refusal awareness. Our results show that the selection-based method demonstrates commendable performance in declining infeasible tasks.
Limitations
Despite the promising results of the proposed Infeasible Benchmark and fine-tuned models, we observe a trade-off between the helpfulness of responses and refusal awareness, suggesting that current approaches are not yet optimal. This identifies a clear avenue for future research. Our current definitions of feasibility are categorized at a coarse level into four groups. Future studies can introduce finer categorizations, which may enable more precise control over the behaviors of LLMs. Given that our research is limited to text-to-text language models, an intriguing direction for future work would be to extend the scope of infeasible task definitions to more advanced models, such as multimodal models or specific AI agents. This expansion could potentially aid in managing and controlling hallucinations more effectively. Additionally, another compelling exploration is to enhance refusal awareness while maintaining the level of helpfulness of these models.
Ethics Statement
Appendix A Benchmark Dataset Summary
Table 6: Summary statistics of Benchmark Dataset.
Feasible Infeasible
n 1850 1898
Length 10.04 9.47
Number for each category 430,531,464,473 1898
Table 7: Summary statistics of long instructions.
Feasible Infeasible
n 145 144
Length 44 60
Appendix B Instruction Tuning Dataset Summary
The dataset we used is under CC-BY-NC-4.0 license.
Table 8: Summary statistics of Instruction Dataset.
Alpaca OOD
# of train split 12784
# of test split 185 200
@@ Appendix B Instruction Tuning Dataset Summary
Appendix C Computing Resources
@@
Appendix D Refusal Expression
@@
Appendix E Prompt Used in Experiments
@@
Figure 8: Prompt for generating feasible data.
Figure 9: Prompt for generating long feasible data.
Figure 10: Prompt for generating long infeasible data.
@@
Figure 15: Prompt for evaluating wint-rate.

### v2

@@ C.1 Long Instruction Benchmark
We have created an additional benchmark dataset focused on long instructions, where each instruction comprises multiple tasks. The dataset is divided into two parts: a feasible subset, where all subinstructions are actionable, and an infeasible subset, which includes a mix of feasible and infeasible subinstructions. An example of an infeasible long instruction is:
To prepare for the upcoming conference, conduct an in-depth literature review on AI trends and compile data from industry reports and academic papers. Develop a detailed presentation, including slides with key statistics and case studies, and attempt to record video lectures summarizing the main points. Gather feedback from the team and attempt to use virtual reality to create an immersive experience for the audience. Coordinate logistics with event organizers, arrange printed materials, and set up a booth for live demonstrations. Post-conference, send thank-you notes, analyze feedback, and prepare a summary report.
@@
The results in Table 6 indicate that long-form instructions are more challenging for current LLMs to accurately determine their feasibility compared to short-form benchmarks. For instance, GPT-4 using the pre-method achieved an AUROC of only 0.865, significantly lower than the 0.965 achieved in the previous short-form benchmark. Also, the overall calibration of probability becomes less well-aligned, which might make the model outputs less trustworthy. Those results highlight the increased difficulty of processing long-form instructions.
C.2 Test Finetuned Models on Alpagasus
@@
We also tested the fine-tuned model on another test dataset called Alpagasus (Chen et al., 2023), which has a large sample size. This dataset contains $700+$ data, carefully curated from multiple resources, and is regarded as "feasible" to LLMs. Since the models we fine-tuned were trained using the Alpaca dataset, we consider this scenario as an evaluation of their ability to handle out-of-distribution data.
@@
The results shown in Table 7 suggest that our fine-tuning method is not very resilient to distribution shifts and may not significantly improve general helpfulness. Therefore, future work should focus on developing more effective instruction-tuning methods to better manage distribution shifts.
Appendix D Computing Resources
@@
Appendix E Refusal Expression
@@
Appendix F Prompt Used in Experiments
@@
Figure 11: Prompt for generating feasible.
@@
Figure 16: Prompt for evaluating win-rate.